OpenStudy (anonymous):

A radio wave has a frequency of 5.5 × 10Ê hertz and travels at a speed of 3.0 × 10§ meters/second. What is its wavelength?

OpenStudy (shamim):

the formula is\[v=f \lambda \]

OpenStudy (shamim):

\[\lambda =\frac{ v }{ f }\]

OpenStudy (priths):

First of all \[\lambda =v/f\] So we simply have to put the values and calculate the answer. So wavelength = v/f